OpenStudy (anonymous):

is x+1=1 always sometimes or never true

OpenStudy (mlgpro):

sometimes

OpenStudy (anonymous):

how so? just trying to understand this for summer school, didn't think it would haunt me.

OpenStudy (mlgpro):

x=2 2+1 doesnt = 1

OpenStudy (asevilla5):

its not true

OpenStudy (asevilla5):

You can think of this in terms of money. If I have x dollars, and you have x + 1 dollars, how could we ever have the same amount of money? (If I have 5 dollars, and you have 5+ 1 = 6 dollars, or if I have 100 dollars and you have 100+1 = 101 dollars, etc.) You will always have one dollar more. Algebraically we start with x + 0 = x + 1 ignore the plus zero x = x + 1 subtract x from both sides 0 = 1 ... um. Zero does not equal one. :) So the statement is NEVER true.

OpenStudy (mlgpro):

go to http://www.mathpapa.com/algebra-calculator.html and type x+1=1

OpenStudy (anonymous):

cool thanks helped a bunch (y)

OpenStudy (mlgpro):

yw